OpenStudy (anonymous):

Alex made a sketch for a poster contest. The sketch measures 4 in.×5 in. The scale factor from the sketch to the finished poster was 2 in. to 5 in. What were the dimensions of the poster?

OpenStudy (anonymous):

10 x 12 1/2

OpenStudy (anonymous):

1.6 in. × 2 in. 6 in. × 10 in. 9 in. × 7 in. 10 in. × 12.5 in.

OpenStudy (anonymous):

The linear dimensions of the poster are 5/2 those of the sketch. 4*5/2 = 10 5*5/2 = 25/2 = 12/5 The poster is 10 x 12.5 inches